OpenStudy (anonymous):

what are the first five terms of the sequence given by the formula a^n=5^n+1?

OpenStudy (anonymous):

a. 1, 5,9, 13,17 b. 6, 11, 16, 21, 26 c. 5, 10, 15, 20, 25 d. 5,6,7,8,9

OpenStudy (anonymous):

b

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

a^n=5^n+1 and question is \[a _{n}=5n+1\] according to your options.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

not a^n=5^n+1

OpenStudy (anonymous):

so simply put n=1,,2,3,4,5 you will get answer

OpenStudy (anonymous):

haha duh lol im sorry

OpenStudy (anonymous):

can you help with another?

OpenStudy (anonymous):

ask as a new question if i can't so then someone else can do.

OpenStudy (anonymous):

okay